# Robust Instance Optimal Phase-Only Compressed Sensing

### v2

@@ A.3 Fixed-$\mathbf{x}$ Bound on the Orthogonal Term
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ $\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}\right)$
$\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}+\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}\right)$
$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\underbrace{\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}}_{:=f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})}\Big{|},$ (A.19)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}$ $\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{*}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}+1=O(1),$
which holds with probability at least $1-4\exp(-cm)$ under $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see (2.10).
It remains to bound $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$, where $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ is defined in (A.19). We first bound $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ for a fixed $\mathbf{x}$ by using the following extended matrix deviation inequality, along with the rotational invariance of $\bm{\Phi}$.
@@
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C[\omega(\mathcal{U}_{c}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2}).$
@@
By $\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}\in\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathcal{U}^{\rm(N)}=\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathbf{P}^{\top}\mathbf{P}\mathcal{U}^{\rm(N)}=(\mathbf{I}_{n}-\mathbf{e}_{1}\mathbf{e}_{1}^{\top})\mathbf{P}\mathcal{U}^{\rm(N)}$, we know the first entry of $\tilde{\mathbf{u}}:=\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}$ is always 00. We further observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$ implies
$\begin{bmatrix}\tilde{\mathbf{u}}\\ \mathbf{w}\end{bmatrix}=\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ \mathbf{w}\end{bmatrix}\in\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\mathcal{U}_{c}^{\rm(N)}:=\tilde{\mathcal{U}}_{0}.$
We let $\tilde{\mathbf{u}}_{1}\in\mathbb{R}^{n-1}$ be the restriction of $\tilde{\mathbf{u}}\in\mathbb{R}^{n}$ to the last $n-1$ entries, and let $\tilde{\mathcal{U}}\subset\mathbb{R}^{m+n-1}$ be the restriction of $\tilde{\mathcal{U}}_{0}\subset\mathbb{R}^{m+n}$ to the last $m+n-1$ entries, so that we have $(\tilde{\mathbf{u}}_{1}^{\top},\mathbf{w}^{\top})^{\top}\in\tilde{\mathcal{U}}$. This can be more precisely formulated as
@@
In light of [49, Exercise 7.5.4] we have $\omega(\tilde{\mathcal{U}})\leq\omega(\tilde{\mathcal{U}}_{0})\leq\omega(\mathcal{U}_{c}^{\rm(N)})$.
With these preparations, we let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}^{[1]},\tilde{\bm{\Phi}}^{[2:n]}]$ where $\tilde{\bm{\Phi}}^{[1]}$ is the first column of $\tilde{\bm{\Phi}}$ and $\tilde{\bm{\Phi}}^{[2:n]}\in\mathbb{R}^{m\times(n-1)}$ is composed by the last $n-1$ columns of $\tilde{\bm{\Phi}}$, and can proceed as follows:
$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.20)
$\qquad=\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}\mathbf{e}_{1}))^{*}\tilde{\bm{\Phi}})\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}$
@@
where $\hat{\bm{\Phi}}:=\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}^{[1]}))^{*}\tilde{\bm{\Phi}}^{[2:n]})$ has the same distribution as a matrix with i.i.d. $\mathcal{N}(0,1)$ entries almost surely. Recalling $\omega(\tilde{\mathcal{U}})\leq\omega(\mathcal{U}_{c}^{\rm(N)})$ and observing $\operatorname{rad}(\tilde{\mathcal{U}})\leq\operatorname{rad}(\tilde{\mathcal{U}}_{0})\leq\operatorname{rad}(\mathcal{U}_{c}^{\rm(N)})\leq 1$, a straightforward application of Lemma A.4 yields the desired claim. ∎
@@
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C_{1}[\omega(\mathcal{U}_{1}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2})-4\exp(-c_{2}m).$ (A.22)
@@
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{3}[\omega(\mathcal{U}_{c}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2})-4\exp(-c_{4}m).$ (A.23)
By (A.5), $m=\Omega(\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\rm(N)}))$, and letting $t=c_{5}\sqrt{m}\delta$ with sufficiently small $c_{5}$ in (A.22)–(A.23), we obtain
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\delta\right)\geq 1-11\exp(-c_{6}\delta^{2}m).$
@@ A.5 Uniform (All-$\mathbf{x}$) Bound on the Parallel Term
Under the setting of Lemma A.3, let $\mathcal{K}$ be an arbitrary subset of $\mathbb{S}^{n-1}$. There exist absolute constants $c_{1},C_{2},C_{3},C_{4}$ such that for any $\eta\in(0,c_{1})$ and $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$, if
$m\geq\frac{C_{2}}{\log(\eta^{-1})}\left(\frac{\omega^{2}(\mathcal{U}_{1}^{\rm(N)})}{\eta^{2}}+\frac{\mathscr{H}(\mathcal{K},r)}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}}\right),$ (A.24)
then with probability at least $1-C_{3}\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$, we have
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq C_{4}\eta\sqrt{\log(\eta^{-1})}.$
@@
$\mathbbm{P}\left(\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq\frac{C(\omega(\mathcal{U}_{1}^{\rm(N)})+t)}{\sqrt{m}}\right)\geq 1-2\exp\big{(}\mathscr{H}(\mathcal{K},r)-t^{2}\big{)}$
for any $t\geq 0$. Therefore, setting $t=\Theta(\sqrt{\mathscr{H}(\mathcal{K},r)}+\omega(\mathcal{U}_{1}^{\rm(N)}))$ yields that the event
$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq\frac{C_{1}[\omega(\mathcal{U}_{1}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},r)}]}{\sqrt{m}}$ (A.25)
holds with probability at least $1-2\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$.
Bounding the Number of Small Measurements: For small enough $\eta>0$, we recall $\mathcal{J}_{\mathbf{x},\eta}:=\big{\{}i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\big{\}}.$ We now bound $|\mathcal{J}_{\mathbf{x},\eta}|$ over $\mathbf{x}\in\mathcal{N}_{r}$. For a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$, by $\Re(\bm{\Phi}_{i}^{*}\mathbf{x})\sim\mathcal{N}(0,1)$, we have $p_{0}:=\mathbbm{P}\big{(}|\bm{\Phi}_{i}\mathbf{x}|\leq\eta\big{)}\leq\mathbbm{P}\big{(}|\Re(\bm{\Phi}_{i}^{*}\mathbf{x})|\leq\eta\big{)}\leq\sqrt{\frac{2}{\pi}}\eta.$ Note that $|\mathcal{J}_{\mathbf{x},\eta}|\sim\text{Bin}(m,p_{0})$, so the Chernoff bound (e.g., [36, Sec. 4.1]) gives $\mathbbm{P}\big{(}|\mathcal{J}_{\mathbf{x},\eta}|\geq\eta m\big{)}\leq\exp(-c_{1}\eta m)$ for some absolute constant $c_{1}$. Thus, a union bound over $\mathbf{x}\in\mathcal{N}_{r}$ gives

### v3

@@ A.3 Fixed-$\mathbf{x}$ Bound on the Orthogonal Term
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$
$\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big|\right)$
$\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}+\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big|\right)$
$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\underbrace{\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}}_{:=f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})}\Big|,$ (A.19)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}$ $\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{*}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}+1=O(1),$
which holds with probability at least $1-4\exp(-cm)$ under $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}}))$; see (2.10).
It remains to bound $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$, where $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ is defined in (A.19). We first bound $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ for a fixed $\mathbf{x}$ by using the following extended matrix deviation inequality, along with the rotational invariance of $\bm{\Phi}$.
@@
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C[\omega(\mathcal{U}_{c}^{\mathbb{S}})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2}).$
@@
By $\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}\in\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathcal{U}^{\mathbb{S}}=\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathbf{P}^{\top}\mathbf{P}\mathcal{U}^{\mathbb{S}}=(\mathbf{I}_{n}-\mathbf{e}_{1}\mathbf{e}_{1}^{\top})\mathbf{P}\mathcal{U}^{\mathbb{S}}$, we know the first entry of $\tilde{\mathbf{u}}:=\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}$ is always 0. We further observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}$ implies
$\begin{bmatrix}\tilde{\mathbf{u}}\\ \mathbf{w}\end{bmatrix}=\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ \mathbf{w}\end{bmatrix}\in\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\mathcal{U}_{c}^{\mathbb{S}}:=\tilde{\mathcal{U}}_{0}.$
We let $\tilde{\mathbf{u}}_{1}\in\mathbb{R}^{n-1}$ be the restriction of $\tilde{\mathbf{u}}\in\mathbb{R}^{n}$ to the last n1n-1 entries, and let $\tilde{\mathcal{U}}\subset\mathbb{R}^{m+n-1}$ be the restriction of $\tilde{\mathcal{U}}_{0}\subset\mathbb{R}^{m+n}$ to the last m+n1m+n-1 entries, so that we have $(\tilde{\mathbf{u}}_{1}^{\top},\mathbf{w}^{\top})^{\top}\in\tilde{\mathcal{U}}$. This can be more precisely formulated as
@@
In light of [50, Exercise 7.5.4] we have $\omega(\tilde{\mathcal{U}})\leq\omega(\tilde{\mathcal{U}}_{0})\leq\omega(\mathcal{U}_{c}^{\mathbb{S}})$.
With these preparations, we let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}^{[1]},\tilde{\bm{\Phi}}^{[2:n]}]$ where $\tilde{\bm{\Phi}}^{[1]}$ is the first column of $\tilde{\bm{\Phi}}$ and $\tilde{\bm{\Phi}}^{[2:n]}\in\mathbb{R}^{m\times(n-1)}$ is composed by the last n1n-1 columns of $\tilde{\bm{\Phi}}$, and can proceed as follows:
$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.20)
$\qquad=\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}\mathbf{e}_{1}))^{*}\tilde{\bm{\Phi}})\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big|$
@@
where $\hat{\bm{\Phi}}:=\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}^{[1]}))^{*}\tilde{\bm{\Phi}}^{[2:n]})$ has the same distribution as a matrix with i.i.d. $\mathcal{N}(0,1)$ entries almost surely. Recalling $\omega(\tilde{\mathcal{U}})\leq\omega(\mathcal{U}_{c}^{\mathbb{S}})$ and observing $\operatorname{rad}(\tilde{\mathcal{U}})\leq\operatorname{rad}(\tilde{\mathcal{U}}_{0})\leq\operatorname{rad}(\mathcal{U}_{c}^{\mathbb{S}})\leq 1$, a straightforward application of Lemma A.4 yields the desired claim. ∎
@@
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C_{1}[\omega(\mathcal{U}_{1}^{\mathbb{S}})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2})-4\exp(-c_{2}m).$ (A.22)
@@
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{3}[\omega(\mathcal{U}_{c}^{\mathbb{S}})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2})-4\exp(-c_{4}m).$ (A.23)
By (A.5), $m=\Omega(\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}}))$, and letting $t=c_{5}\sqrt{m}\delta$ with sufficiently small $c_{5}$ in (A.22)–(A.23), we obtain
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\delta\right)\geq 1-11\exp(-c_{6}\delta^{2}m).$
@@ A.5 Uniform (All-$\mathbf{x}$) Bound on the Parallel Term
Under the setting of Lemma A.3, let $\mathcal{K}$ be an arbitrary subset of $\mathbb{S}^{n-1}$. There exist absolute constants $c_{1},C_{2},C_{3},c_{4},C_{5}$ such that for any $\eta\in(0,c_{1})$ and $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$, if
$m\geq\frac{C_{2}}{\log(\eta^{-1})}\left(\frac{\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}})}{\eta^{2}}+\frac{\mathscr{H}(\mathcal{K},r)}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}}\right),$ (A.24)
then with probability at least $1-C_{3}\exp(-c_{4}\eta^{2}\log(\eta^{-1})m)$, we have
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|\leq C_{5}\eta\sqrt{\log(\eta^{-1})}.$
@@
$\mathbbm{P}\left(\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|\leq\frac{C(\omega(\mathcal{U}_{1}^{\mathbb{S}})+t)}{\sqrt{m}}\right)\geq 1-2\exp\big(\mathscr{H}(\mathcal{K},r)-t^{2}\big)$
for any $t\geq 0$. Under the sample complexity $m=\Omega(\frac{1}{\eta^{2}\log(\eta^{-1})}[\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}})+\mathscr{H}(\mathcal{K},r)])$, setting $t=\eta\sqrt{\log(\eta^{-1})}\sqrt{m}$ yields that the event
$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|\leq C_{1}\eta\sqrt{\log(\eta^{-1})}$ (A.25)
holds with probability at least $1-2\exp(-\frac{1}{2}\eta^{2}\log(\eta^{-1})m)$.
Bounding the Number of Small Measurements: For small enough $\eta>0$, we recall $\mathcal{J}_{\mathbf{x},\eta}:=\big\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\big\}.$ We now bound $|\mathcal{J}_{\mathbf{x},\eta}|$ over $\mathbf{x}\in\mathcal{N}_{r}$. For a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$, by $\Re(\bm{\Phi}_{i}^{*}\mathbf{x})\sim\mathcal{N}(0,1)$, we have $p_{0}:=\mathbbm{P}\big(|\bm{\Phi}_{i}\mathbf{x}|\leq\eta\big)\leq\mathbbm{P}\big(|\Re(\bm{\Phi}_{i}^{*}\mathbf{x})|\leq\eta\big)\leq\sqrt{\frac{2}{\pi}}\eta.$ Note that $|\mathcal{J}_{\mathbf{x},\eta}|\sim\text{Bin}(m,p_{0})$, so the Chernoff bound (e.g., [37, Sec. 4.1]) gives $\mathbbm{P}\big(|\mathcal{J}_{\mathbf{x},\eta}|\geq\eta m\big)\leq\exp(-c_{1}\eta m)$ for some absolute constant $c_{1}$. Thus, a union bound over $\mathbf{x}\in\mathcal{N}_{r}$ gives